OpenStudy (anonymous):

What is f(-2) for f(x)=-1/2x+5?

OpenStudy (anonymous):

plug -2 into equation as x also careful when you enter in equations here we dont know if you mean -1/(2x+5) or -1/2x +5

OpenStudy (anonymous):

ok, but how do i solve it. will it look like f(-2)=-1/2x+5

OpenStudy (anonymous):

it will look like f(-2) = - 1/2(-2) + 5

OpenStudy (anonymous):

1/4 + 5 = 5.25

OpenStudy (anonymous):

thanks! :)